Circular No. 6608

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

NOVALIKE VARIABLE IN SAGITTARIUS S. Kimeswenger, H. Gratl, and F. Kerber, University of Innsbruck;

P. Fouque, Observatoire de Paris, Meudon; S. Kohle, Radioastronomisches Institut, Universität Bonn; and S. Steele, University of Toronto Southern Observatory (UTSO), write: “We have followed the evolution of of Sakurai’s object (IAUC 6322, 6601) since Apr. 1996. Infrared-imaging photometry (IJK) obtained on 1997 Mar. 24.25 UT at the ESO 1-m telescope (+ DeNIS infrared survey instrument) shows that the flux has increased by 100 percent in I, 315 percent in J, and 940 percent in K (I = 9.29, J = 7.86, K = 6.24), as compared to earlier observations (see Kerber et al. 1997, IAU Symp. 180; Duerbeck and Benetti (1996, Ap.J. 468, L111). Visual (BV RI_{c}) CCD photometry obtained on Mar. 28.29–28.33 with the Dutch 0.90-m telescope at La Silla and the 0.60-m UTSO telescope at Las Campanas shows that the object has faded by ∼ 40 percent in B, is constant at V, and has increased 75 and 100 percent in R and I, respectively (B = 12.58, V = 11.17, R = 9.96, I = 9.26). The observed spectral distribution cannot be explained by a single stellar component. Using the color temperature of ∼ 5000 K obtained from the visual data, we find that the near-infrared excess (of about a factor of 4 at 2.1 μm) can be explained by a blackbody of ∼ 1800 K producing a specific output 30 times larger than the stellar source, evidently the signature of hot, extended, circumstellar dust.”

LMC X-4 D. Dal Fiume, F. Frontera, and M. Orlandini, TESRE, Bologna; A.

Santangelo, S. Del Sordo, IFCAI, Palermo; L. Piro, IAS, Frascati; A. Coletta, G. Celidonio, G. D’Andreta, G. Gandolfi, G. Spoliti, and D. Ricci, BeppoSAX, Rome; and J. M. Muller, SRON, Utrecht, report: “The BeppoSAX Wide Field Cameras detected five major flares from the x-ray pulsar LMC X-4 during Mar. 21.7882–21.7981 UT. The most prominent flare reached a flux level of ∼ 300 mCrab (2–26 keV). The duration of each flare was ∼ 100 s. The pre-/post-flaring flux level of LMC X-4 is estimated at ∼ 28 mCrab (2–26 keV). The x-ray emission from LMC X-4 during flares is remarkably softer than in the pre-/post-flaring phases. Our detection is consistent both in shape and intensity with the earlier GINGA detection (Levine et al. 1991, Ap.J. 381, 101). There is no appreciable γ-ray flux increase corresponding to the observed flares, with a 90-percent upper limit of 110 mCrab above 40 keV.”

1997 March 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green